# A modern perspective on rational homotopy theory

### v1

@@ Abstract.
In Quillen’s paper on rational homotopy theory, the category of $1$-reduced simplicial sets is endowed with a family of model structures, the most prominent of which is the one whose weak equivalences are the rational homotopy equivalences. In this paper, we give a modern approach to this family of model structures.
@@ 1. Introduction
Rational homotopy theory studies $1$-connected spaces by simplifying the computation of their homotopy groups. The non-vanishing homotopy groups of a $1$-connected space are abelian, so their torsion can be discarded by tensoring with the rational numbers. Discarding torsion makes homotopy group computations more tractable: in the case of spheres, their rational homotopy groups were computed by Serre in [Ser51]. The success of rational homotopy theory has led to its extension to nilpotent spaces by Bousfield and Kan in [BK72]; Sullivan in [Sul74] and [Sul77]; Hilton, Mislin, and Roitberg in [HMR75]; Bousfield and Gugenheim in [BG76]; and Neisendorfer in [Nei78]. Ultimately, rational homotopy theory has been extended to all spaces by Casacuberta and Peschke in [CP93]; Gómez-Tato, Halperin, and Tanré in [GTHT00]; and Bastardas and Casacuberta in [BC01].
Model categories, as introduced by Quillen in [Qui67], provide a formal framework in which many tools and constructions from homotopy theory are defined. For instance, every model category admits a homotopy category, which is the localization of the model category at its class of weak equivalences. Thus, a model structure on a category endows the category with a homotopy theory and gives us tools to study that homotopy theory.
@@
In [Qui69], Quillen applies his recently introduced notion of a model category to study rational homotopy theory, and especially how rational homotopy theory translates difficult problems in geometry and topology to more feasible problems in algebra. More specifically, Quillen produces model categories both on the topological side and on the algebraic side, and then proves that specific functorial constructions between them induce equivalences of these model categories. The notion of equivalence that we are referring to is that of a Quillen equivalence; in particular, a Quillen equivalence induces an equivalence of homotopy categories. On the algebraic side, Quillen produces model structures on rational differential graded (DG) Lie algebras and rational DG coalgebras, which are important algebraic invariants for computations in rational homotopy theory.
@@
In this paper, we focus on the topological side, where Quillen produces a family of model structures on the category of $1$-reduced simplicial sets: simplicial sets whose $1$-skeleton is a point. As the category of $1$-connected spaces is ill-behaved with respect to important types of limits and colimits, such as pullbacks, the passage to $1$-reduced simplicial sets is necessary for Quillen’s model-categorical approach to rational homotopy theory.
@@
This family of model structures interpolates between ordinary homotopy theory, where the weak equivalences are weak homotopy equivalences, and rational homotopy theory, where the weak equivalences are rational homotopy equivalences. Namely, for a multiplicative subset $M$ of $\mathbb{Z}$, Quillen endows the category of $1$-reduced simplicial sets with a model structure in which a map is a weak equivalence if and only if it induces an isomorphism of homotopy groups after tensoring with $M^{-1}\mathbb{Z}$; we call such a map an $M$-local homotopy equivalence.
Tensoring with $M^{-1}\mathbb{Z}$ for different choices of a multiplicative subset $M$ of $\mathbb{Z}$ allows us to preserve specific types of torsion in homotopy groups. For instance, for a prime $p$, taking $M=\mathbb{Z}\setminus p\mathbb{Z}$ and tensoring with the $p$-local integers $\mathbb{Z}_{(p)}$ specifically preserves $p$-torsion. Thus, we can study the torsion one prime at a time.
@@
$M$$\{1\}$ $\mathbb{Z}\setminus 2\mathbb{Z}$$\mathbb{Z}\setminus\{0\}$
[…]
torsion discarded by $-\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$ none all but $2$-torsion all
$M$-local homotopy equivalence weak $2$-local rational
To establish this family of model structures, Quillen verifies the model category axioms one-by-one. In later years, though, further advancements in the study of model categories have provided us with powerful tools that enable us to induce new model structures from old without needing to perform such a painstaking verification.
@@
The motivating idea behind our work in this paper is that we can use these modern tools to provide a new approach to Quillen’s family of model structures on $1$-reduced simplicial sets. The first tool we employ is that of left transfer, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [BHK+15]. This tool allows us to transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. Then, we left Bousfield localize the latter model structure to obtain the rest of Quillen’s family. Although the tool of left Bousfield localization dates back to the work of Bousfield in [Bou75] and [Bou77] and Farjoun in [Far92] and [Far95b], we rely on the more recent existence results for left Bousfield localization developed by Hirschhorn in [Hir03] and by Barwick in [Bar10].
@@
For the left transfer, we observe that a $1$-reduced simplicial set has a unique 00-simplex, so there is a forgetful functor from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets. This functor admits a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set as defined by May in [May67]; in fact, Quillen uses this construction to establish his family of model structures. As the $1$st Eilenberg subcomplex depends on the basepoint, the passage to pointed simplicial sets, rather than unpointed ones, is necessary. We use this adjunction to left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to the following model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$.
Theorem 1.1.
There is a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
@@
    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
@@
    a cofibration if and only if $f$ is a monomorphism.
Then, for the left Bousfield localization, we characterize, for a multiplicative subset $M$ of $\mathbb{Z}$, the $M$-local spaces and the $M$-local homotopy equivalences in the language of function complexes. Our characterizations can be viewed as the enriched-in-simplicial-sets version of the work of Casacuberta, Peschke, and Pfenniger in [CPP92], which uses hom-sets rather than function complexes. Using these characterizations, we left Bousfield localize the model category from Theorem 1.1 to recover the entire family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$ from [Qui69]. We say that a simplicial set $X$ is $M$-local if $X$ is a Kan complex whose higher homotopy groups are $M^{-1}\mathbb{Z}$-modules.
Theorem 1.2.
@@
    a cofibration if and only if $f$ is a monomorphism.
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local.
We also give an alternative description of this family of model structures using Bousfield localization with respect to homology, as introduced in [Bou75], in conjunction with the result on left-induced model structures from [BHK+15] that we use for Theorem 1.1. The idea is that, rather than first left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localize, we can try to localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homology with $M^{-1}\mathbb{Z}$-coefficients and then left-transfer this localized model structure, which we denote by $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$, to ${\mathcal{SS}\mathrm{ets}}_{1}$. We show that this approach also yields the family of model structures from Theorem 1.2. Schematically, the diagram of model-categorical operations
[…]
commutes, thus giving us two modern descriptions of the family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$.
Another motivating goal of ours is to show that these model structures have properties that one generally desires and expects in the context of model categories. The tools that we invoke guarantee us most of these properties. An exception is the property of being a simplicial model category, as defined by Quillen in [Qui67]. We establish this property explicitly by adapting the techniques of Goerss and Jardine in [GJ99] to show that the model category from Theorem 1.1 is simplicial. As left Bousfield localization preserves the simplicial structure, this structure is bestowed upon the entire family of model structures on $1$-reduced simplicial sets from Theorem 1.2. We remark that, in [Bur21], Burke proves the analog of our simplicial structure for reduced simplicial sets: simplicial sets with a unique 00-simplex. However, Burke’s line of argument is different from that in this paper.
Also, we partially extend Theorem 1.2 to the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets. Namely, we produce a family of model structures on simplicial sets that describes a generalization of rational homotopy theory to all spaces. In [Iva22], Ivanov gives an overview of the various approaches to rational homotopy theory beyond the framework of $1$-connectedness or nilpotency. We adopt the approach of Gómez-Tato, Halperin, and Tanré from [GTHT00]. In this approach, one rationalizes a connected space $X$ by applying fiberwise rationalization to the fibration
@@
given by the $1$st Postnikov approximation of $X$; its homotopy fiber $\widetilde{X}$ is the universal cover of $X$, which is a $1$-connected space that can be rationalized in the classical sense. The tool of fiberwise localization is invoked by Sullivan in [Sul74] and investigated by Llerena in [Lle82] and [Lle85] and Farjoun in [Far95a]. This rationalization theory for non-nilpotent spaces is also studied in recent work of Rivera, Wierstra, and Zeinalian in [RWZ21].
We partially extend Theorem 1.2 to all simplicial sets as follows.
Theorem 1.3.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which a simplicial set $X$ is fibrant if and only if $X$ is $M$-local.
We also prove the counterpart of this theorem for the category ${\mathcal{T}\mathrm{op}}$ of topological spaces. The reason why Theorem 1.3 only partially extends Theorem 1.2 to all simplicial sets is that we lack a description of the weak equivalences other than the standard one for a left Bousfield localization. This deficiency is due to the fact that fiberwise rationalization satisfies a more intricate universal property than that of the rationalization of $1$-connected spaces, which causes our characterization of the weak equivalences in Theorem 1.2 to break down.
@@
Lastly, we verify that our localization $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ of spaces with respect to rational homotopy from Theorem 1.3 does not coincide with Bousfield’s localization of spaces with respect to rational homology from [Bou75]. This result captures the fact that rational homotopy and rational homology do not agree beyond the framework of $1$-connected spaces; in fact, our proof of the result involves a specific instance of this general phenomenon.
Organization of the paper

### v2

@@ Abstract.
In Quillen’s paper on rational homotopy theory, the category of $1$-reduced simplicial sets is endowed with a family of model structures, the most prominent of which is the one in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational Kan complexes. In this paper, we give a modern approach to this family of model structures. We recover Quillen’s family of model structures by first left-transferring the model structure on pointed simplicial sets and then left Bousfield localizing at the rationalization maps of spheres. Applying this localization to the model category of all spaces yields a model category in which the weak equivalences are the rational homotopy equivalences in the extended sense of Gómez-Tato, Halperin, and Tanré and the fibrant objects are the rational spaces. Thus, we generalize Quillen’s family of model structures beyond the rational homotopy theory of $1$-connected spaces.
Key words and phrases:
rational homotopy theory, model categories, left-induced model structures, left Bousfield localization.
1991 Mathematics Subject Classification:
55P62 (Primary), 55P60, 55U10, 55U35, 18N55, 18N50, 18N40 (Secondary).
@@ 1. Introduction
Classically, rational homotopy theory is the study of $1$-connected spaces up to the torsion in their homotopy groups. The non-vanishing homotopy groups of a $1$-connected space are abelian, so their torsion can be discarded by tensoring with the rational numbers. In this framework, we adjust our notion of equivalence from that of a weak homotopy equivalence, which preserves all homotopy groups on the nose, to that of a rational homotopy equivalence, which induces isomorphisms of higher homotopy groups only after tensoring with the rational numbers. Conceptually, a rational homotopy equivalence can be thought of as a weak homotopy equivalence up to an error in torsion. In the $1$-connected setting, one can equivalently use rational homology instead; a map of $1$-connected spaces is a rational homotopy equivalence if and only if it is a rational homology isomorphism.
Accordingly, the corresponding notion of a nice space in rational homotopy theory is a space whose higher homotopy groups are already rational vector spaces, that is, uniquely divisible by every positive integer; such spaces are called rational. In [39], Sullivan constructs, for a $1$-connected space $X$, a relative CW complex $(X_{\mathbb{Q}},X)$ such that $X_{\mathbb{Q}}$ is a rational space and the relative CW inclusion is a rational homotopy equivalence. This construction is called the rationalization of $X$, and it identifies the rational homotopy groups of $X$ with the higher homotopy groups of $X_{\mathbb{Q}}$ in a topological fashion.
The geometric construction that underpins the rationalization of spaces is the rationalization of the $n$-sphere $S^{n}$ in dimensions $n\geq 2$ and, more specifically, the construction of the rational $n$-sphere $S^{n}_{\mathbb{Q}}$. In essence, the rationalization of a general space is built out of the CW inclusions of $n$-spheres into rational $n$-spheres, hence the resulting relative CW construction. Thus, the rationalization maps of spheres are the building blocks for the rational homotopy theory of spaces.
Working rationally can simplify computations significantly. In the case of spheres, though their homotopy groups are notoriously difficult to compute, their rational homotopy groups were computed by Serre in [36]. This contrast captures the fact that the difficulty in computing the homotopy groups of spheres is concentrated in their torsion. It also suggests that, while a direct computation of homotopy groups can be too ambitious an approach, the computation of localizations of homotopy groups can be a more tractable goal.
1.1. Quillen’s approach to rational homotopy theory
The computational features of rational homotopy theory extend deeper than the systematic disregard of torsion in homotopy groups. Rational homotopy theory can be translated purely into algebraic terms, and rational computations in geometry and topology can be transformed into more feasible ones in algebra. Quillen’s paper on rational homotopy theory [33] establishes this advantage of rational homotopy theory, and it informs us about the algebra that fully encodes the rational homotopy theory of $1$-connected spaces. Specifically, Quillen’s work turns the study of the rational homotopy theory of $1$-connected spaces into that of rational differential graded Lie algebras and rational differential graded coalgebras up to quasi-isomorphism.
Quillen’s approach to rational homotopy theory takes place within his framework of model categories from [32]. A model structure on a category allows one to introduce tools and constructions from homotopy theory in that category, and thus endows the category with a homotopy theory. Specifically, a model structure on a category comprises a class of morphisms called weak equivalences together with two other classes of morphisms, cofibrations and fibrations, subject to a list of axioms. This structure allows for constructions such as the homotopy category of a model category, which is its localization at the class of weak equivalences. In practice, model categories often satisfy additional properties that render the study of their homotopy theory less abstract and more explicit.
In his paper on rational homotopy theory, Quillen produces a number of model structures, including
  1. (1)
    one on $1$-connected spaces whose weak equivalences are the rational homotopy equivalences;
  2. (2)
    one on rational differential graded Lie algebras whose weak equivalences are the quasi-isomorphisms; and
  3. (3)
    one on rational differential graded coalgebras whose weak equivalences are also the quasi-isomorphisms,
@@ 1.1. Quillen’s approach to rational homotopy theory
as well as functorial constructions that induce equivalences of these model categories. The notion of equivalence that we are referring to is, in fact, called a Quillen equivalence. Thus, Quillen’s work quantifies how rational homotopy theory translates difficult problems in geometry and topology into more feasible problems in algebra. In other words, Quillen uses his framework of model categories to model the rational homotopy theory of $1$-connected spaces with the homotopy theory of rational differential graded Lie algebras and coalgebras.
On the side of $1$-connected spaces, Quillen does not work so much with topological spaces as with simplicial sets. Simplicial sets can be thought of as simplicial complexes with degeneracy data that allows us to view lower-dimensional simplices as honorary higher-dimensional simplices, as well as directionality data for all simplices. This point of view illustrates that simplicial sets are a combinatorial model for spaces, and the degeneracy and directionality data render them a more faithful model than that of ordinary simplicial complexes. An exposition of this approach to simplicial sets and their advantages over simplicial complexes is provided by Bergner in [6].
@@
Instead of working directly with $1$-connected topological spaces, Quillen produces a family of model structures on the category of $1$-reduced simplicial sets, that is, simplicial sets whose $1$-skeleton is a point. As the category of $1$-connected topological spaces is ill-behaved with respect to important limits and colimits, such as pullbacks, the passage to $1$-reduced simplicial sets is necessary for Quillen’s model-categorical approach, which requires a category that has all finite limits and colimits. Thus, Quillen uses $1$-reduced simplicial sets as a combinatorial analog of $1$-connected spaces in analogy with how simplicial sets are a combinatorial model for spaces, except that, now, the more rigid structure of a $1$-reduced simplicial set yields a nicer underlying category to work in than that of $1$-connected spaces.
More specifically, Quillen produces a family of model structures on $1$-reduced simplicial sets that interpolates between ordinary homotopy theory, whose weak equivalences are the weak homotopy equivalences, and rational homotopy theory, whose weak equivalences are the rational homotopy equivalences. This interpolation captures the fact that we have options between discarding no torsion at all, as is the case with weak homotopy equivalences, and discarding all torsion, as is the case with rational homotopy equivalences. For instance, given a prime number $p$, we can specifically preserve $p$-torsion in higher homotopy groups by tensoring them with the $p$-local integers $\mathbb{Z}_{(p)}$, and we can adapt our notion of equivalence to be that of a $p$-local homotopy equivalence, which induces isomorphisms of higher homotopy groups only after tensoring with the $p$-local integers. Such a selective approach allows us to study the torsion in homotopy groups one prime at a time rather than study it in one go or discard it altogether.
@@
In full detail, for a multiplicative subset $M$ of $\mathbb{Z}$, Quillen endows the category of $1$-reduced simplicial sets with a model structure in which a map is a weak equivalence if and only if it induces isomorphisms of higher homotopy groups after tensoring with $M^{-1}\mathbb{Z}$; we call such a map an $M$-local homotopy equivalence. Tensoring with $M^{-1}\mathbb{Z}$ for different choices of $M$ allows us to preserve specific types of torsion in homotopy groups. For instance, for a prime number $p$, choosing $M=\mathbb{Z}\setminus p\mathbb{Z}$ and tensoring with $\mathbb{Z}_{(p)}$ only keeps $p$-torsion.
In this notation, a space is called $M$-local if its higher homotopy groups are $M^{-1}\mathbb{Z}$-modules, that is, uniquely divisible by every positive integer in $M$. Moreover, for a $1$-connected space $X$, Sullivan’s geometric construction yields a relative CW complex $(X_{M^{-1}\mathbb{Z}},X)$ such that $X_{M^{-1}\mathbb{Z}}$ is an $M$-local space and the relative CW inclusion is an $M$-local homotopy equivalence. This construction is called the $M$-localization of $X$. Thus, rational homotopy theory can be developed for any localization of $\mathbb{Z}$, though specific computations, such as Serre’s computation of the rational homotopy groups of spheres, much depend on the choice of such a localization.
@@
$M$ $\{1\}$ $\mathbb{Z}\setminus p\mathbb{Z}$ $\mathbb{Z}\setminus\{0\}$
$M^{-1}\mathbb{Z}$ $\mathbb{Z}$ $\mathbb{Z}_{(p)}$ $\mathbb{Q}$
torsion discarded by $\otimes M^{-1}\mathbb{Z}$ none all but $p$-torsion all
$M$-local homotopy equivalence weak $p$-local rational
1.2. A modern perspective on Quillen’s approach
@@
To establish his family of model structures on $1$-reduced simplicial sets, Quillen painstakingly verifies the model category axioms one-by-one. The motivating idea behind our work in this paper is that we can use modern tools to provide a streamlined approach to this family of model structures. The first tool we employ is that of left transfer, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [4]. This tool allows us to transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. Then, we left Bousfield localize the latter model structure at the $M$-localization maps of spheres to recover the rest of Quillen’s family. Although the tool of left Bousfield localization dates back to the work of Bousfield [7, 8] and Farjoun [19, 21], we rely on the more recent existence results for left Bousfield localization developed by Hirschhorn in [26] and by Barwick in [2].
@@
For the left transfer, we observe that a $1$-reduced simplicial set has a unique 0-simplex, so there is a forgetful functor from the category of $1$-reduced simplicial sets to the category of pointed simplicial sets. This functor admits a right adjoint given by the 11st Eilenberg subcomplex of a pointed simplicial set, as defined by May in [30]; in fact, Quillen uses this construction to give his family of model structures. We use the 11st Eilenberg subcomplex functor to left-transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences and whose fibrant objects are the Kan complexes.
Then, for the left Bousfield localization, we characterize, for a multiplicative subset $M$ of $\mathbb{Z}$, the $M$-local spaces and the $M$-local homotopy equivalences using function complexes. Our characterizations can be viewed as the enriched-in-simplicial-sets version of the work of Casacuberta, Peschke, and Pfenniger in [15], which involves hom-sets instead. Using these characterizations, we left Bousfield localize our left-induced model category at the $M$-localization maps of spheres to obtain a model structure on $1$-reduced simplicial sets whose weak equivalences are the $M$-local homotopy equivalences and whose fibrant objects are the $M$-local Kan complexes. Thus, we recover the entirety of Quillen’s family of model structures on $1$-reduced simplicial sets.
In particular, we obtain a model structure on the category of $1$-reduced simplicial sets in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational Kan complexes. This model category encapsulates the rational homotopy theory of $1$-connected spaces; for instance, fibrant replacements in this model category are rationalizations of $1$-connected spaces. It is this model category that Quillen establishes to be equivalent to his model structures on rational differential graded Lie algebras and rational differential graded coalgebras.
We also give an alternative approach using Bousfield localization with respect to homology, as introduced by Bousfield in [7], in conjunction with left transfer. The idea is that, for a multiplicative subset $M$ of $\mathbb{Z}$, rather than first left-transfer the model structure on pointed simplicial sets and then localize, we can try to localize the model structure on pointed simplicial sets with respect to homology with $M^{-1}\mathbb{Z}$-coefficients and then left-transfer this localized model structure. We show that this approach also recovers Quillen’s family of model structures on $1$-reduced simplicial sets.
Schematically, for a multiplicative subset $M$ of $\mathbb{Z}$, we denote
@@
    the model category of pointed simplicial sets by $\mathcal{SS}\mathrm{ets}$$_{\ast}$};
@@
    the model structure on $1$-reduced simplicial sets with weak homotopy equivalences by $\mathcal{SS}\mathrm{ets}_{1}$};
@@
    the localization of $\mathcal{SS}\mathrm{ets}$$_{\ast}$} with respect to homology with $M^{-1}\mathbb{Z}$-coefficients by $\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$}; and
@@
    the model structure on $1$-reduced simplicial sets with $M$-local homotopy equivalences by $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{1}$},
and we show that the diagram of model-categorical operations
$\mathcal{SS}\mathrm{ets}$$_{\ast}$$\mathcal{SS}\mathrm{ets}_{1}$$\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$$\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{1}$}}localizationtransferlocalizationtransfer
commutes. As the weak equivalences in $\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$} are the $M$-local homology isomorphisms, the commutativity of the diagram stems from the result that a map of $1$-connected spaces is an $M$-local homotopy equivalence if and only if it is an $M$-local homology isomorphism.
Another motivating goal of ours is to show that these model structures on $1$-reduced simplicial sets have properties that one generally desires and expects in the context of model categories. The tools that we invoke guarantee us most of these properties. An exception is the property of being a simplicial model category, as defined by Quillen in [32]. We establish this property explicitly by first showing that our left-induced model structure on $1$-reduced simplicial sets is simplicial. As left Bousfield localization preserves the simplicial structure, this structure is bestowed upon the entire family of model structures on $1$-reduced simplicial sets.
1.3. Generalizing Quillen’s family of model structures to all spaces
The modern tools that we use to recover Quillen’s family of model structures on $1$-reduced simplicial sets allow us to extend it to all spaces. This extension encodes the generalization of rational homotopy theory to non-simply connected spaces, which has been accomplished in two stages. Firstly, rational homotopy theory was extended to nilpotent spaces by Sullivan [37, 38, 39]; Bousfield and Kan [11]; Hilton, Mislin, and Roitberg [25]; Bousfield and Gugenheim [10]; and Neisendorfer [31]. Ultimately, it was extended to all spaces by Casacuberta and Peschke [14]; Gómez-Tato, Halperin, and Tanré [24]; and Bastardas and Casacuberta [3]. At both stages of generalization, rational homotopy theory has been translated to purely algebraic terms, as in the classical case.
We use our modern perspective to produce a family of model structures on the category of spaces that describes the generalization of rational homotopy theory to non-simply connected spaces due to Gómez-Tato, Halperin, and Tanré. In their approach, one rationalizes a connected space $X$ by applying fiberwise rationalization to the fibration
@@ 1.3. Generalizing Quillen’s family of model structures to all spaces
given by the 11st Postnikov approximation of $X$, where $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space. The fiber $\widetilde{X}$ is the universal cover of $X$, which is a $1$-connected space that can be rationalized in the classical sense. This technique generalizes to disconnected spaces by fiberwise rationalizing every component. The tool of fiberwise localization is invoked by Sullivan [37] and investigated by Llerena [28, 29] and Farjoun [20], and the specific case of fiberwise rationalization is treated in recent work of Ivanov [27].
In general, for a multiplicative subset $M$ of $\mathbb{Z}$, the fiberwise $M$-localization of a connected space $X$ entails, at the level of total spaces, a map $X\rightarrow X_{M^{-1}\mathbb{Z}}$ that induces isomorphisms on $\pi_{0}$ and $\pi_{1}$ and isomorphisms of higher homotopy groups after tensoring with $M^{-1}\mathbb{Z}$. We call such a map an $M$-local homotopy equivalence, for it extends the notion of an $M$-local homotopy equivalence beyond maps of $1$-connected spaces. In [34], Rivera, Wierstra, and Zeinalian give algebraic characterizations for rational homotopy equivalences in this general sense, which they refer to as $\pi_{1}$-rational homotopy equivalences. Also, as in the $1$-connected case, a space is called $M$-local if its higher homotopy groups are $M^{-1}\mathbb{Z}$-modules; note that no conditions are imposed on $\pi_{0}$ or $\pi_{1}$.
In this terminology, we endow the category of spaces with a model structure whose weak equivalences are the $M$-local homotopy equivalences and whose fibrant objects are the $M$-local spaces. This model category is the left Bousfield localization of the model category of spaces at the $M$-localization maps of spheres. In particular, we get a model structure on the category of spaces in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational spaces. In [24], Gómez-Tato, Halperin, and Tanré produce an algebraic category with a notion of homotopy between its morphisms whose homotopy category is equivalent to the homotopy category of rational spaces in their extended sense. By our theorem, the latter homotopy category arises from a model category, so it can be studied using the structured toolkit of model categories.
@@
Lastly, we note that our localization of all spaces at rational homotopy does not coincide with Bousfield’s localization of spaces at rational homology; the unique map from $\mathbb{R}P^{2}$ to the point is a rational homology isomorphism, but not a rational homotopy equivalence in the extended sense, for it does not induce an isomorphism of fundamental groups. This result captures the fact that rational homotopy and rational homology do not agree for non-simply connected spaces.
1.4. Organization of the paper